# Coarse-Grained Boltzmann Generators

### v1

@@ Abstract
Sampling equilibrium molecular configurations from the Boltzmann distribution is a longstanding challenge. Boltzmann Generators (BGs) address this by combining exact-likelihood generative models with importance sampling, but their practical scalability is limited. Meanwhile, coarse-grained surrogates enable the modeling of larger systems by reducing effective dimensionality, yet often lack the reweighting process required to ensure asymptotically correct statistics. In this work, we propose Coarse-Grained Boltzmann Generators (CG-BGs), a principled framework that unifies scalable reduced-order modeling with the exactness of importance sampling. CG-BGs act in a coarse-grained coordinate space, using a learned potential of mean force (PMF) to reweight samples generated by a flow-based model. Crucially, we show that this PMF can be efficiently learned from rapidly converged data via force matching. Our results demonstrate that CG-BGs faithfully capture complex interactions mediated by explicit solvent within highly reduced representations, establishing a scalable pathway for the unbiased sampling of larger molecular systems.
@@ 1 Introduction
Accurately sampling molecular configurations from the Boltzmann distribution is a fundamental challenge in statistical physics (Chandler, 1987). Access to these samples is essential for computing thermodynamic observables, including free energies (Chipot and Pohorille, 2007). For molecular systems, the high dimensionality renders direct evaluation of the partition function intractable, forcing a reliance on simulation methods such as Molecular Dynamics (MD) or Markov Chain Monte Carlo (MCMC) (Frenkel and Smit, 2002). However, these approaches often struggle with the rugged, multi-modal structure of potential energy landscapes. High free-energy barriers trap trajectories in metastable states, leading to highly correlated samples and slow convergence (Lindorff-Larsen et al., 2011). Despite extensive progress in enhanced sampling methods (Zhu et al., 2025; Hénin et al., 2022) (e.g., umbrella sampling (Torrie and Valleau, 1977), metadynamics (Laio and Gervasio, 2008)) and coarse-graining (Noid, 2013; Saunders and Voth, 2013), efficient equilibrium sampling remains a major bottleneck.
@@
Deep generative models have recently emerged as a promising alternative for equilibrium sampling (Noé et al., 2019; Albergo et al., 2019; Wirnsberger et al., 2020). A notable example is the Boltzmann Generator (BG) (Noé et al., 2019). By learning a diffeomorphic transformation between a simple prior (e.g., a Gaussian) and the complex molecular configuration space, BGs enable efficient proposal generation and exact likelihood evaluation. The tractable likelihood permits rigorous reweighting of generated configurations to the target Boltzmann distribution via importance sampling, yielding asymptotically exact equilibrium statistics. This enables amortized sampling (Tan et al., 2025b), where generation cost is substantially lower than MD simulations.
@@
Figure 1: CG-BG workflow. (1) Training data are collected and mapped from atomistic configurations to CG beads. (2) A PMF network learns $U_{\eta}(\mathbf{R})$ from rapidly converged data, while a normalizing flow learns a proposal density $q_{\theta}(\mathbf{R})$. (3) CG samples from flow models are reweighted with the PMF to recover the target distribution $p(\mathbf{R})$ and compute unbiased thermodynamic observables.
@@
In practice, however, scaling BGs to high-dimensional molecular systems remains challenging (Klein and Noé, 2024; Tan et al., 2025a). As system size increases, even expressive generative models exhibit diminishing overlap with the target distribution, resulting in high-variance importance weights and ineffective reweighting. Moreover, exact likelihood evaluation requires computing Jacobian determinants (Chen et al., 2018), which incurs significant computational overhead and scales poorly with dimensionality (Hutchinson, 1989). As a result, current BGs have mostly been applied to small systems, such as short peptides (Tan et al., 2025b; Rehman et al., 2025), often restricted to empirical implicit solvent models (Hawkins et al., 1995; Nguyen et al., 2013).
@@
Coarse-graining (CG) provides a complementary approach by projecting atomistic configurations onto a lower-dimensional set of collective variables. This idea underlies Boltzmann Emulators (Lewis et al., 2025; Zheng et al., 2024; Jing et al., 2024; Zhu et al., 2026) and related generative surrogates (Daigavane et al., 2025; Schreiner et al., 2023; Plainer et al., 2025; Costa et al., 2025; dos Santos Costa et al., 2024; Diez et al., 2025; Vlachas et al., 2021; Xu et al., 2025), which reduce the number of degrees of freedom and allow models to effectively handle larger systems. However, these approaches face a circular dependency: training generally requires unbiased equilibrium data, which is precisely the computational bottleneck they aim to overcome. In most cases, CG surrogates are often trained on finite-time, non-equilibrium trajectories (Lewis et al., 2025; Zheng et al., 2024) and, unlike Boltzmann Generators, most lack a reweighting procedure to correct for this mismatch, resulting in biased statistics.
@@
Present work. In this paper, we introduce Coarse-Grained Boltzmann Generators (CG-BGs, Fig. 1), a class of Boltzmann Generators that operate directly in CG coordinate  [^1] Our code is publicly available at https://github.com/tummfm/cg-bg.. CG-BGs combine generative modeling with importance sampling using a learned potential of mean force (PMF) as the target energy, enabling exact equilibrium sampling in a reduced-dimensional representation. This design provides a scalable pathway for sampling high-dimensional molecular systems. CG-BGs are particularly advantageous as they can be trained directly from rapidly converged data and effectively capture complex solvent-mediated effects.
@@
  • We introduce Coarse-Grained Boltzmann Generators, a scalable framework for equilibrium sampling in CG coordinate space. To our knowledge, this is the first BG formulation that incorporates machine learning potentials (MLPs) as the target energy for importance sampling.
  • We leverage enhanced sampling force matching to learn CG PMFs directly from rapidly converged simulations, eliminating reliance on unbiased equilibrium data and providing a correction mechanism for CG Boltzmann Emulators.
  • We demonstrate that CG-BGs capture solvent-mediated interactions in highly reduced representations, achieving improved accuracy over classical implicit solvent models while dramatically reducing computational cost relative to atomistic BGs.
@@ 2 Background and Preliminaries
where $\beta=(k_{B}T)^{-1}$ and $Z$ is the partition function. The goal of equilibrium sampling is to generate samples from $p(\mathbf{r})$ in order to compute observables $\mathbb{E}_{p}[\mathcal{O}]=\int\mathcal{O}(\mathbf{r})p(\mathbf{r})d\mathbf{r}$, and free energies. A dataset is considered converged if the empirical averages of observables statistically recover equilibrium expectations.
@@ 2.1 Boltzmann Generators and Emulators
Boltzmann Generators (BGs) (Noé et al., 2019) combine exact-likelihood generative models with importance sampling to obtain exact equilibrium statistics. Typically implemented using normalizing flows (Rezende and Mohamed, 2016), a BG defines a tractable proposal density $q_{\theta}(\mathbf{r})$ that approximates $p(\mathbf{r})$. Given samples $\mathbf{r}_{i}\sim q_{\theta}$, importance weights are assigned as
@@
Provided that $q_{\theta}$ overlaps sufficiently with $p$, this procedure converges to the exact Boltzmann average. This allows BGs to be trained on biased or non-equilibrium data, since reweighting restores equilibrium statistics at inference time.
Boltzmann Emulators adopt similar generative architectures but omit the reweighting step, estimating observables directly under $q_{\theta}$. While computationally efficient at inference, emulators are no longer asymptotically exact: their accuracy depends entirely on how closely $q_{\theta}$ approximates $p(\mathbf{r})$. This dependence necessitates high-quality unbiased training data, whose generation is often the primary computational bottleneck.
@@ 2.2 Continuous Normalizing Flows
Continuous normalizing flows (CNFs) extend discrete normalizing flows (Rezende and Mohamed, 2016; Dinh et al., 2014) by modeling density transformations as solutions to time-dependent ordinary differential equations (ODEs) (Chen et al., 2018). A vector field $v_{\theta}:[0,1]\times\mathbb{R}^{n}\to\mathbb{R}^{n}$, parameterized by a neural network, defines the dynamics
@@
where $p_{0}$ is a simple prior distribution. The solution at time $t$ is given by
@@
While CNFs are commonly trained via maximum likelihood, Flow Matching (FM) (Albergo et al., 2023; Lipman et al., 2022; Liu et al., 2022) provides a simulation-free alternative. Conditional Flow Matching (CFM) (Tong et al., 2023) directly regresses the neural vector field $v_{\theta}$ onto a target conditional vector field $u_{t}(\mathbf{x}\mid z)$ that induces a prescribed probability path. The training objective is
@@
where $t\sim\mathcal{U}[0,1]$ and $z$ denotes a conditioning variable. A commonly used construction specifies the probability path through linear interpolation between paired samples (Albergo et al., 2023; Tong et al., 2023). Specifically, let $z=(\mathbf{x}_{0},\mathbf{x}_{1})$, where $\mathbf{x}_{0}$ and $\mathbf{x}_{1}$ are samples drawn from the source and target distributions, respectively. The interpolated state and corresponding vector field are
@@ 2.3 Coarse-Graining and Potentials of Mean Force
Coarse-graining reduces dimensionality by mapping atomistic configurations $\mathbf{r}\in\mathbb{R}^{n}$ to a lower-dimensional set of collective variables (CVs), or beads, $\mathbf{R}\in\mathbb{R}^{N}$ ($N\ll n$), typically selected to capture the slow degrees of freedom, via a mapping $\mathbf{R}=\Xi(\mathbf{r})$. From a bottom-up perspective (Noid et al., 2008; Jin et al., 2022), the goal of coarse-graining is to define a suitable set of CG variables and parameterize an effective potential such that the CG model reproduces the marginal equilibrium distribution of the underlying atomistic system:
@@
The PMF captures both energetic and entropic contributions from the eliminated degrees of freedom and generally induces many-body, state-dependent interactions (Krishna et al., 2009).
In practice, the PMF is intractable due to the high-dimensional integral over $\mathbf{r}$. Classical CG force fields (Marrink et al., 2007; Souza et al., 2021) approximate $U(\mathbf{R})$ using fixed functional forms, which may lack sufficient expressivity. Recent approaches parameterize $U(\mathbf{R})$ using neural networks trained via variational principles such as force matching (Noid et al., 2008; Wang et al., 2019) or relative entropy minimization (Shell, 2008; Thaler et al., 2022).
@@ 3 Coarse-Grained Boltzmann Generators
Atomistic BGs provide exact equilibrium sampling but scale poorly to large systems, while CG Boltzmann Emulators improve scalability at the cost of persistent bias due to the absence of reweighting.
We propose CG-BGs that reconcile scalability with statistical rigor by performing generative modeling and importance sampling directly in CG coordinate space. Rather than targeting the full atomistic distribution, CG-BGs target the marginal distribution $p(\mathbf{R})$ defined by the PMF.
CG-BGs have two major components: a flow-based generative model to propose CG samples (similar to atomistic BGs) and a machine learning PMF acting as an energy predictor for importance sampling. Unlike atomistic MLPs trained on potential energy surfaces (Behler and Parrinello, 2007; Blank et al., 1995), CG MLPs have no ground-truth energy labels, since the PMF is a free energy that includes entropic contributions from eliminated degrees of freedom. We next describe how the PMF is learned from atomistic simulations and outline the CG-BG workflow (Fig. 1).
@@ 3.1 Variational Force Matching
The central premise is that the CG forces predicted by the model should match, on average, the instantaneous atomistic forces projected onto the CG coordinates, denoted as $\mathcal{F}_{\mathrm{proj}}(\mathbf{r})$. The gradient of the exact PMF satisfies:
@@
where the expectation is taken over the fiber distribution (Hummerich et al., 2025), i.e., the conditional probability of a fine-grained configuration $\mathbf{r}$ given its CG representative $\mathbf{R}$.
From a learning perspective, instantaneous projected forces provide noisy samples to approximate the target mean force. Given a dataset $\mathcal{D}$ of atomistic configurations, a parameterized CG potential $U_{\eta}(\mathbf{R})$ is trained by minimizing
@@
While global LSI conditions are strong assumptions for multimodal PMFs (Vempala and Wibisono, 2019), this proposition theoretically motivates minimizing force errors as a proxy for distributional accuracy. This is important for the subsequent importance sampling step, as the quality of the reweighted ensemble depends directly on the divergence between the learned target and the true marginal.
@@ 3.2 Enhanced Sampling for Force Matching
Standard force matching relies on unbiased equilibrium data, which can be prohibitively expensive to obtain for systems with metastable states and large free-energy barriers—a limitation shared by Boltzmann Emulators. Moreover, the Boltzmann distribution decays exponentially with energy, leading to sparse sampling of high-energy transition regions that are critical for accurately reconstructing the global PMF.
Enhanced sampling force matching (ESFM) (Chen et al., 2026) overcomes these limitations by exploiting the invariance of the fiber (conditional) distribution under coarse-grained biasing.
{restatable} propositionfiber (Chen et al. (2026)) Let $V(\mathbf{R})$ be a bias potential depending only on the coarse-grained coordinates. The conditional distribution of atomistic configurations given $\mathbf{R}$ remains unchanged: $p_{V}(\mathbf{r}\mid\mathbf{R})=p(\mathbf{r}\mid\mathbf{R}).$ (15)
Since the mean force $\nabla U(\mathbf{R})$ is the expectation of the projected forces over this conditional distribution (Eq. 12), the force regression targets remain unbiased even when sampled from a biased equilibrium ensemble. ESFM therefore minimizes
@@
where $\mathcal{D}_{\mathrm{bias}}$ is a rapidly converged dataset generated via enhanced sampling and $\mathcal{F}_{\mathrm{proj}}(\mathbf{r})$ denotes forces recomputed using the unbiased atomistic potential.
@@
Together, Propositions 3.2 (Proof in §A.2) and 3.2 establish that ESFM enables accurate PMF learning from rapidly converged data, benefiting from faster convergence and improved coverage of transition regions.

### v2

@@ Abstract
Sampling equilibrium molecular configurations from the Boltzmann distribution is a longstanding challenge. Boltzmann Generators (BGs) address this by combining exact-likelihood generative models with importance sampling, but practical scalability is limited. Meanwhile, coarse-grained surrogates enable the modeling of larger systems by reducing effective dimensionality, yet often lack a reweighting procedure required to ensure asymptotically correct statistics. In this work, we propose Coarse-Grained Boltzmann Generators (CG-BGs), a framework for reduced-order generative modeling with importance sampling in coarse-grained coordinate space. CG-BGs generate samples using a flow-based model and reweight them using a learned potential of mean force (PMF). We show that the PMF can be learned from rapidly converged trajectories via enhanced sampling force matching. Experiments demonstrate that CG-BGs capture solvent-mediated interactions in highly reduced representations while substantially reducing computational cost relative to atomistic BGs, providing a practical route toward equilibrium sampling of larger molecular systems.
@@ 1 Introduction
Accurately sampling molecular configurations from the Boltzmann distribution is a central problem in statistical physics (Chandler, 1987). These samples are required for estimating observables and thermodynamic quantities, including free energies (Chipot and Pohorille, 2007). For molecular systems, the high dimensionality of the configuration space makes direct computation of the partition function intractable, forcing a reliance on simulation methods such as Molecular Dynamics (MD) or Markov Chain Monte Carlo (MCMC) (Frenkel and Smit, 2002). However, these methods are often inefficient in systems with rugged energy landscapes. High free energy barriers lead to metastable trapping, producing strongly correlated samples and slow convergence (Lindorff-Larsen et al., 2011). Despite extensive progress in enhanced sampling methods (Zhu et al., 2025; Hénin et al., 2022) (e.g., umbrella sampling (Torrie and Valleau, 1977), metadynamics (Laio and Gervasio, 2008)) and coarse-graining (Noid, 2013; Saunders and Voth, 2013), equilibrium sampling at scale remains difficult.
@@
Deep generative models have recently emerged as a promising alternative for equilibrium sampling (Noé et al., 2019; Albergo et al., 2019; Wirnsberger et al., 2020). A notable example is the Boltzmann Generator (BG) (Noé et al., 2019). By learning a diffeomorphic transformation between a simple prior (e.g., a Gaussian) and the complex molecular configuration space, BGs enable efficient proposal generation and exact-likelihood evaluation. The tractable likelihood permits rigorous reweighting of generated configurations to the target Boltzmann distribution via importance sampling, allowing unbiased estimation of equilibrium observables. This has enabled amortized sampling strategies (Tan et al., 2025b), where generation is substantially cheaper than MD simulations.
@@
Figure 1: CG-BG framework. Atomistic configurations are mapped to CG coordinates to construct training data. A PMF network learns $U_{\eta}(\mathbf{R})$ from rapidly converged data, while a normalizing flow learns a proposal density $q_{\theta}(\mathbf{R})$ over CG configurations. Samples generated from the flow are reweighted using the PMF to recover the target equilibrium distribution $p(\mathbf{R})$, enabling unbiased estimation of thermodynamic observables.
@@
In practice, however, scaling BGs to high-dimensional molecular systems remains difficult (Klein and Noé, 2024; Tan et al., 2025a). As system size increases, even expressive generative models exhibit diminishing overlap with the target distribution, resulting in high-variance importance weights and ineffective reweighting. In addition, likelihood evaluation requires Jacobian determinant computation (Chen et al., 2018), which introduces significant computational overhead and scales poorly with dimensionality (Hutchinson, 1989). As a result, existing BG applications are largely restricted to small systems such as short peptides (Tan et al., 2025b; Rehman et al., 2025) with empirical implicit solvent models (Hawkins et al., 1995; Nguyen et al., 2013).
@@
Coarse-graining (CG) provides a complementary approach by projecting atomistic configurations onto a lower-dimensional set of collective variables. This idea underlies Boltzmann Emulators (Lewis et al., 2025; Zheng et al., 2024; Jing et al., 2024; Zhu et al., 2026) and related generative surrogates (Schreiner et al., 2023; Daigavane et al., 2025; Plainer et al., 2025; Costa et al., 2025; dos Santos Costa et al., 2024; Diez et al., 2025; Vlachas et al., 2021; Xu et al., 2025). By reducing the number of degrees of freedom during generation, these methods can be applied to larger systems. However, training requires converged unbiased simulation data, which is difficult to obtain in practice. Thus, CG models are often trained on short finite-time trajectories that do not fully capture the target distribution (Lewis et al., 2025; Zheng et al., 2024) and, unlike Boltzmann Generators, do not incorporate reweighting to correct this mismatch due to the lack of an explicit energy function for the target distribution, resulting in biased statistical estimates.
@@
Present work. In this paper, we introduce Coarse-Grained Boltzmann Generators (CG-BGs, Fig. 1), a class of Boltzmann Generators that operate directly in CG coordinates  [^1] Our code is publicly available at https://github.com/tummfm/cg-bg.. CG-BGs combine generative modeling with importance sampling using a learned potential of mean force (PMF) as the target energy, enabling asymptotically correct equilibrium sampling in a reduced-dimensional representation. This design provides a scalable pathway for sampling high-dimensional molecular systems. CG-BGs are particularly advantageous as they can be trained directly from rapidly converged data and effectively capture complex solvent-mediated effects.
@@
  • We introduce Coarse-Grained Boltzmann Generators, a scalable framework for equilibrium sampling in CG coordinate space using machine learning potentials (MLPs) as the target energy for importance sampling.
  • We show that enhanced sampling force matching enables learning the PMF from rapidly converged simulation trajectories, eliminating reliance on unbiased equilibrium data and providing a correction mechanism for Boltzmann Emulators.
  • We demonstrate that CG-BGs capture solvent-mediated interactions in highly reduced representations, achieving improved accuracy over classical implicit solvent models while substantially reducing computational cost relative to atomistic BGs.
@@ 2 Background and Preliminaries
where $\beta=(k_{B}T)^{-1}$ and $Z$ is the partition function. The goal of equilibrium sampling is to generate samples from $p(\mathbf{r})$ in order to compute observables $\mathbb{E}_{p}[\mathcal{O}]=\int\mathcal{O}(\mathbf{r})p(\mathbf{r})d\mathbf{r}$, and free energies. A dataset is considered converged if empirical averages of observables match their equilibrium expectations within statistical error.
@@ 2.1 Boltzmann Generators and Emulators
Boltzmann Generators (BGs) (Noé et al., 2019) combine exact-likelihood generative models with importance sampling to estimate equilibrium properties. Typically implemented using normalizing flows (Rezende and Mohamed, 2016), a BG defines a tractable proposal density $q_{\theta}(\mathbf{r})$ that approximates $p(\mathbf{r})$. Given samples $\mathbf{r}_{i}\sim q_{\theta}$, importance weights are computed as
@@
Provided that $q_{\theta}$ overlaps sufficiently with $p$, this estimator converges to the corresponding Boltzmann averages. This allows BGs to be trained on biased or non-equilibrium samples, with reweighting correcting the induced distribution shift at evaluation.
Boltzmann Emulators adopt similar generative architectures but omit the reweighting step, relying directly on $q_{\theta}$ for estimating observables. Model accuracy is therefore determined by the quality of the learned distribution $q_{\theta}$, with no correction applied at inference time. This places stronger requirements on the training data: accurate models require long unbiased trajectories, which are difficult to obtain in practice.
@@ 2.2 Continuous Normalizing Flows
Continuous normalizing flows (CNFs) extend discrete normalizing flows (Dinh et al., 2014; Rezende and Mohamed, 2016) by modeling density transformations as solutions to time-dependent ordinary differential equations (ODEs) (Chen et al., 2018). A vector field $v_{\theta}:[0,1]\times\mathbb{R}^{n}\to\mathbb{R}^{n}$, parameterized by a neural network, defines the dynamics
@@
where $p_{0}$ is a simple prior distribution. The solution at time $t$ is
@@
While CNFs are often trained using maximum likelihood, Flow Matching (FM) (Lipman et al., 2022; Liu et al., 2022; Albergo et al., 2023) provides a simulation-free alternative. Conditional Flow Matching (CFM) (Tong et al., 2023) directly regresses the neural vector field $v_{\theta}$ onto a target conditional vector field $u_{t}(\mathbf{x}\mid z)$ that induces a prescribed probability path. The training objective is
@@
where $t\sim\mathcal{U}[0,1]$ and $z$ is a conditioning variable. A common choice uses linear interpolation between paired source and target samples (Albergo et al., 2023; Tong et al., 2023). Let $z=(\mathbf{x}_{0},\mathbf{x}_{1})$ with $\mathbf{x}_{0}$ and $\mathbf{x}_{1}$ sampled from the source and target distributions, respectively. The interpolated state and corresponding vector field are
@@ 2.3 Coarse-Graining and Potentials of Mean Force
Coarse-graining maps atomistic configurations $\mathbf{r}\in\mathbb{R}^{n}$ to a lower-dimensional set of collective variables (CVs), or beads, $\mathbf{R}\in\mathbb{R}^{N}$ with $N\ll n$, through a mapping $\mathbf{R}=\Xi(\mathbf{r})$. The CG variables are typically chosen to retain the slow degrees of freedom. In bottom-up coarse-graining (Noid et al., 2008; Jin et al., 2022), the objective is to construct an effective CG potential such that the CG model reproduces the marginal equilibrium distribution of the atomistic system:
@@
The PMF includes both energetic and entropic contributions from the eliminated degrees of freedom and generally contains many-body, state-dependent interactions (Krishna et al., 2009).
Evaluating the PMF is intractable in practice due to the high-dimensional integral over $\mathbf{r}$. Classical CG force fields (Marrink et al., 2007; Souza et al., 2021) approximate $U(\mathbf{R})$ using fixed functional forms, which may lack sufficient expressivity. More recent approaches represent $U(\mathbf{R})$ using neural networks trained by force matching (Noid et al., 2008; Wang et al., 2019) or relative entropy minimization (Shell, 2008; Thaler et al., 2022).
@@ 3 Coarse-Grained Boltzmann Generators
Atomistic BGs permit reweighting-based equilibrium sampling but become difficult to apply to large systems. Boltzmann Emulators improve scalability by omitting reweighting, at the cost of introducing bias.
We introduce CG-BGs, which perform generative modeling and importance sampling directly in CG coordinate space. Instead of the full atomistic distribution, CG-BGs target the marginal distribution $p(\mathbf{R})$ defined by the PMF.
CG-BGs consist of two components: a flow-based model that generates CG configurations and a learned PMF used for importance reweighting. Unlike atomistic MLPs trained on labeled energies (Blank et al., 1995; Behler and Parrinello, 2007), CG PMFs cannot be directly evaluated from atomistic configurations because they include entropic contributions from eliminated degrees of freedom. We next describe how the PMF is learned from atomistic simulations and outline the CG-BG workflow (Fig. 1).
@@ 3.1 Variational Force Matching
The central condition is that CG forces should match, in expectation, the instantaneous atomistic forces projected onto the CG coordinates, denoted by $\mathcal{F}_{\mathrm{proj}}(\mathbf{r})$. The exact PMF satisfies:
@@
where the expectation is taken over the fiber distribution (Hummerich et al., 2025), i.e., the conditional distribution of atomistic configuration $\mathbf{r}$ given $\mathbf{R}$.
From a learning perspective, instantaneous projected forces provide stochastic estimates of the conditional mean force in Eq. (12). Given a dataset $\mathcal{D}$ of atomistic configurations, a parameterized CG potential $U_{\eta}(\mathbf{R})$ is trained by minimizing
@@
While global LSI conditions are strong assumptions for multimodal PMFs (Vempala and Wibisono, 2019), this result motivates force matching as a proxy for distributional accuracy. This is relevant for importance sampling, where performance depends on the divergence between the learned and true marginal distributions.
@@ 3.2 Enhanced Sampling for Force Matching
Standard force matching requires unbiased converged data, which is expensive to obtain for systems with metastable states and large free energy barriers—a limitation shared by Boltzmann Emulators. In addition, high-energy transition regions are rarely visited under the Boltzmann distribution, yet are important for accurately learning the PMF.
Enhanced sampling force matching (ESFM) (Chen et al., 2026) overcomes these limitations by using invariance of the fiber distribution under coarse-grained biasing.
{restatable} propositionfiber (Chen et al. (2026)) Let $V(\mathbf{R})$ be a bias potential depending only on the coarse-grained coordinates. The conditional distribution of atomistic configurations given $\mathbf{R}$ is invariant: $p_{V}(\mathbf{r}\mid\mathbf{R})=p(\mathbf{r}\mid\mathbf{R}).$ (15)
Since the mean force $\nabla U(\mathbf{R})$ is an expectation of the projected forces over this conditional distribution (Eq. 12), the regression target is unchanged under CG biasing. ESFM minimizes
@@
where $\mathcal{D}_{\mathrm{bias}}$ is a rapidly converged dataset generated using enhanced sampling and $\mathcal{F}_{\mathrm{proj}}(\mathbf{r})$ denotes forces recomputed from the unbiased atomistic potential.
@@
Together, Propositions 3.2 (Proof in §A.2) and 3.2 establish that ESFM enables accurate PMF learning from biased enhanced sampling data, benefiting from faster convergence and improved coverage of transition regions.